Suppose an oven’s radiation wavelength is 0.125 m. A container with 350.00 g of water was placed in the oven, and the temperature of the water rose from 20.0°C to 80.0°C. How many photons of this microwave radiation were required to heat the water? (Assume that all the energy from the radiation was used to raise the temperature of the water.)
Answer #1

2 2 dry het no of phatons - n o No Energy of a photons nhe , h = 6n226X103 / (= 3X108 m/s DT= 80-20= 60c to Speciffe heat of